# HIP-LLM: A Hierarchical Imprecise Probability Approach to Reliability Assessment of Large Language Models

### v1

@@ 5 Discussion
As discussed earlier in Def. 4, the i.i.d. assumption is pragmatically justified in the context of LLM evaluations when tasks are executed in isolation, such as initiating new chat sessions without contextual memory or cross-task influence. Nevertheless, caution is warranted: in practice, the way tasks are designed, sequenced, or batched can introduce subtle dependencies that violate this idealization. Assessors should therefore ensure that testing procedures (e.g., resetting chat sessions) are aligned with the independence assumption.
Indeed, the i.i.d. Bernoulli trial assumption remains a simplification, motivated by mathematical convenience and interpretability. Future work should explicitly incorporate uncertainty or “doubt” about this assumption to examine how sensitive reliability claims are to potential task-level dependencies of outcomes [71].
@@
That said, in this work, our focus is not on prescribing a specific definition of LLM failure but on providing a reliability assessment modeling framework that is agnostic to failure definitions. HIP-LLM can accommodate any well-defined notion of failure, enabling users to instantiate the model with the criteria most relevant to their domain or evaluation objectives.
@@
OPs are fundamental to software reliability assessment, and the same holds true for LLM reliability. While the hierarchical structure of HIP-LLM provides “interfaces” for integrating OPs, readers may nevertheless question the validity of our OP (which are based on benchmark datasets rather than real operational data). Indeed, in the absence of access to real operational data, this study relies on simulated operational data derived from existing benchmark datasets, treating these datasets as sampling frames that approximate the true unknown operational data distribution. Specifically, we assume task-level distributions within each subdomain and assign operational weights across subdomains and domains. We note, as a reliability assessment modeling paper, accurately estimating OPs from real usage data is beyond our current scope. However, this is a well-recognized challenge in software reliability engineering, for which numerous estimation methods have been proposed (e.g., [25, 26, 28]). The concept of OP has also been discussed and measured in AI and ML contexts [24, 76, 78, 79].
The challenge of modeling OPs does not undermine the framework itself but highlight why the OP is essential: it provides the mechanism to reweight and reinterpret benchmark results in terms of real usage. As more representative datasets or user-usage logs [27] become available, the OP can be updated accordingly, making the reliability estimates more faithful to actual field performance.
We acknowledge that our present treatment of OP uncertainty remains simplified: assuming fixed task distributions in each subdomain and constant (sub-)domain weights. More complex modeling on OP uncertainties [72, 73, 26, 74, 75], e.g., dynamic OPs , will form our important future works.
@@ 6 Conclusion
While we believe HIP-LLM represents an important first step toward a principled reliability assessment framework for LLMs (and more broadly, for multimodal generative AI) several key limitations remain (cf. Section 5). These include the challenge of formally defining what constitutes a failure for an LLM, accurately estimating OPs, and systematically eliciting imprecise prior knowledge. Future extensions of this work should aim to address these limitations, for example by developing models for uncertain and dynamically evolving OPs and by integrating more refined mechanisms for prior elicitation.
@@ A.2 Proof of Theorem 1
By substituting Eqs. 5, 6, 7 into 8 we have:
@@ A.3 Proof of Theorem 2
Then, for any $h_{i}\in\mathcal{A}_{i}$, the posterior distribution of $p_{i}$ is characterized by its cumulative distribution function:
@@ A.4 Proof of Theorem 3
Then, for any $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the posterior distribution of $p_{L}$ is characterized by its cumulative distribution function:
@@
We derive the posterior distribution of the LLM reliability $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$ by characterizing its cumulative distribution function, building on the domain-level results from Theorem 2.
@@ A.5 Reliability for Multiple Consecutive Operations
This appendix clarifies the mathematical framework for extending single-operation reliability to $n_{F}$ consecutive operations, which is formalized in Theorems 4, 5, and 6. We explain:
@@
A.5.1 Definition: $n_{F}$-Operation Reliability
@@
    Each operation in subdomain $S_{ij}$ succeeds independently with probability $\theta_{ij}$
@@ A.5.2 Why the CDF Approach
Therefore, we characterize the distribution of $R_{ij}(n_{F})$ through its cumulative distribution function (CDF), which has a tractable integral representation and is sufficient for all practical reliability calculations (e.g., computing probabilities, quantiles, expected values).
@@ A.5.3 Deriving the CDF Formula
$F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})=\Pr(R_{ij}(n_{F})\leq r\mid C_{i},h_{i})=\Pr(\theta_{ij}^{n_{F}}\leq r\mid C_{i},h_{i})$
To compute this probability, we use the following key observation. If $\theta_{ij}^{n_{F}}\leq r$, then taking the $n_{F}$-th root of both sides:
$\theta_{ij}\leq r^{1/n_{F}}$
Since $\theta_{ij}\in(0,1)$ and $r\in[0,1]$, both sides are positive, and the function $x\mapsto x^{1/n_{F}}$ is monotonically increasing on $[0,1]$ for $n_{F}>0$. Therefore, the inequality is preserved under this transformation.
@@
$\Pr(\theta_{ij}\leq r^{1/n_{F}}\mid C_{i},h_{i})=\int_{0}^{r^{1/n_{F}}}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij}$
@@ Appendix C List of Mathematical Notations
$F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})$ CDF of subdomain $n_{F}$-operation reliability
$F_{R_{i}(n_{F})}(r\mid C_{i},h_{i})$ CDF of domain $n_{F}$-operation reliability
$F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H})$ CDF of LLM $n_{F}$-operation reliability

### v2

@@ 5 Discussion
As discussed earlier in Def. 4 and Remark 1, the i.i.d. assumption is pragmatically justified in the context of LLM evaluations when tasks are executed in isolation, such as initiating new chat sessions without contextual memory. Nevertheless, caution is warranted: in practice, the way tasks are designed, sequenced, or batched can introduce subtle dependencies that violate this idealization, e.g., in long-context, agentic workflows. Assessors should therefore ensure testing procedures (e.g., resetting each task session) align with the independence assumption, a common setup in offline benchmarking.
Indeed, the i.i.d. Bernoulli trial assumption remains a simplification, motivated by mathematical convenience and interpretability. However, this assumption is not introduced uniquely by our HIP-LLM, rather that models with the i.i.d. assumption have a long history of being used in software reliability assessment. Thayer et al.’s [67] model was one of the earliest, used in early works on random testing [70]. Regulatory bodies also have recommended using the i.i.d. assumption in reliability assessments when appropriate [3]. Having said that, weakening the i.i.d. assumption has been a long-standing research topic in the software reliability community. Numerous models have been proposed to capture dependence among successive tests, including a binary Markov chain model [17], a Markov renewal process [25, 21], the Markov-based model with benign-failure states [10], and Bayesian models examining how sensitive reliability claims are to potential dependencies of test outcomes [58, 59].
A road-map for extending HIP-LLM beyond i.i.d. task outcomes can build directly on aforementioned, established software reliability models. More concretely, a first step is treating a multi-turn interaction as the unit of analysis and allowing within-session dependence while modeling across-session dynamics via latent states. One can introduce Markov/hidden-state structures, e.g., hierarchical hidden Markov model or Markov renewal formulations where latent “interaction states” evolve over time and govern task-level failure probabilities.
@@
The choice of scoring criteria for defining LLM failures is inherently application dependent and should align with domain-specific specifications (if any) and evaluation norms. For example, in code generation, failures are typically defined in terms of functional correctness, such as compilation success or passing predefined test cases. In mathematical problem solving, correctness is often determined by agreement with a unique ground-truth answer or a formally verifiable solution. In contrast, in legal analysis, failure definitions may rely on expert judgment regarding factual accuracy, consistency with legal principles, or omission of critical considerations. For more open-ended tasks such as email drafting or creative writing, failures are often defined relative to stylistic, pragmatic, or preference-based criteria, frequently assessed through human evaluation. These examples illustrate that failure definitions for LLMs are not universal but must be specified in accordance with the intended use and domain requirements, reinforcing the need to interpret reliability estimates as conditional on the chosen scoring criteria.
That said, in this work, our focus is not on prescribing a specific definition of LLM failure but on providing a reliability assessment modeling framework that is agnostic to failure definitions. For example, the reliability definition in Def. 4 can be readily extended from binary failures to non-binary scores by replacing the indicator function with a continuous scoring function (e.g., taking values in $[0,1]$ through probabilistic or multi-rater failure models), thereby capturing scoring noise or subjectivity. HIP-LLM can accommodate any well-defined notion of failure, enabling users to instantiate the model with the criteria most relevant to their domain or evaluation objectives.
@@
OPs are fundamental to software reliability assessment, and the same principle applies to LLM reliability. HIP-LLM provides explicit hierarchical “interfaces” for integrating OPs; however, in the absence of access to real operational data, this study relies on simulated OPs derived from existing benchmark datasets. Specifically, benchmark datasets are treated as sampling frames that approximate the unknown operational data distribution (cf. Remark 7), with task-level distributions assumed within each subdomain and operational weights assigned across subdomains and domains. This experimental choice is made solely to demonstrate how HIP-LLM incorporates OPs and to enable fair comparison with existing benchmark-based methods, and should not be interpreted as a solution to OP acquisition. Accurately estimating OPs from real usage data is beyond the scope of this reliability modeling study and is a well-recognized, separate research problem in software reliability engineering, for which numerous methodologies have been proposed (e.g., [50, 49, 44]). The concept of OP has also been discussed and measured in AI software contexts [22, 27, 88, 84].
Importantly, the challenge of modeling LLM OPs does not undermine the proposed framework; rather, it highlights why OPs are essential to reliability assessment. In established software reliability practice, OPs are progressively refined as new information becomes available (e.g., execution logs, usage telemetry, or expert-informed adjustments) without requiring changes to the underlying reliability model [50, 49, 44, 62]. As more representative datasets, user interaction logs, or task-distribution estimates for LLMs become available [15], OPs can be updated accordingly, allowing reliability estimates to be recalibrated toward actual field usage. Such separation and “decoupling design” of HIP-LLM between evidence (testing results) and usage assumptions (OPs) enables HIP-LLM to evolve naturally with deployment conditions, improving fidelity without redesigning the model.
While we acknowledge that our present treatment of OP uncertainty remains simplified (for demonstration/comparison), more complex modeling on OP uncertainties, e.g., dynamic OPs, will form our important future works. Notably, a range of reliability models in traditional software engineering explicitly account for OP uncertainty and dynamics, including Bayesian updating, adaptive OP adjustment, and black-box or time-varying usage models [9, 49, 52, 53, 54]. These established techniques provide a foundation that can be retrofitted into HIP-LLM to support uncertain and dynamical OPs in LLM deployments.
@@
On the practicality of HIP-LLM
HIP-LLM is designed to replace current LLM assessment practices that rely primarily on benchmark scores. As highlighted by the 5 gaps identified in the introduction, benchmark-based evaluations often ignore OPs, overlook dependencies between related domains, and cannot incorporate prior knowledge, limiting their relevance to real-world use. In contrast, HIP-LLM produces decision-oriented reliability assessments by integrating OPs, hierarchical domain dependencies, and epistemic uncertainty and priors. Rather than reporting isolated scores (which, strictly speaking, cannot be regarded as “reliability” under standardized definitions such as ANSI [1]), HIP-LLM enables rigorous reliability claims such as “the confidence that a model will complete a specified number of future tasks without failure under a given usage profile”. This allows stakeholders to compare LLMs within specific domains, aggregate reliability across domains, or assess how reliability changes across different user groups or usage patterns. Such capabilities support informed offline decision-making, including model selection, risk assessment, and regulatory certification prior to deployment. While this work focuses on offline assessment, HIP-LLM could potentially be extended to online reliability monitoring with appropriate engineering optimizations for efficiency (e.g., caching and sliding data windows), which we leave for future work.
@@ 6 Conclusion
While we believe HIP-LLM represents an important first step toward a principled reliability assessment framework for LLMs (and more broadly, for multimodal generative AI) several key limitations remain (cf. Section 5). These include the challenge of formally defining what constitutes a failure for an LLM, accurately estimating OPs, and systematically eliciting imprecise prior knowledge. Moreover, the applicability of HIP-LLM is limited to reset or single-task usage scenarios conforming to the i.i.d. assumption (cf. Remark 1); and should not be overgeneralized to long-context or agentic LLM workflows, which require dependency-aware reliability models—an important future work. More future extensions should aim to develop methods for handling uncertain and dynamically evolving OPs, and integrating more refined mechanisms for prior elicitation.
@@ A.2 Proof of Theorem 1
By substituting Eqs. 8, 9, 10 into 11 we have:
@@ A.3 Proof of Theorem 2
Then, for any $h_{i}\in\mathcal{A}_{i}$, the posterior distribution of $p_{i}$ is characterized by its CDF:
@@ A.4 Proof of Theorem 3
Then, for any $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the posterior distribution of $p_{L}$ is characterized by its CDF:
@@
We derive the posterior distribution of the LLM reliability $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$ by characterizing its CDF, building on the domain-level results from Theorem 2.
@@ A.5 Reliability for Multiple Consecutive Operations
This appendix clarifies the mathematical framework for extending the case $n_{F}=1$ to $n_{F}>1$ (reliability with the required number of future failure-free tasks), which is formalized in Theorems 4, 5, and 6. We explain:
@@
A.5.1 Definition: Reliability with the required $n_{F}$ failure-free future tasks
@@
    Each task in subdomain $S_{ij}$ succeeds independently with probability $\theta_{ij}$
@@ A.5.2 Why the CDF Approach
Therefore, we characterize the distribution of $R_{ij}(n_{F})$ through its CDF, which has a tractable integral representation and is sufficient for all practical reliability calculations (e.g., computing probabilities, quantiles, expected values).
@@ A.5.3 Deriving the CDF Formula
$F_{R_{ij}(n_{F})}(t\mid C_{i},h_{i})=\Pr(R_{ij}(n_{F})\leq t\mid C_{i},h_{i})=\Pr(\theta_{ij}^{n_{F}}\leq t\mid C_{i},h_{i})$
To compute this probability, we use the following key observation. If $\theta_{ij}^{n_{F}}\leq t$, then taking the $n_{F}$-th root of both sides:
$\theta_{ij}\leq t^{1/n_{F}}$
Since $\theta_{ij}\in(0,1)$ and $t\in[0,1]$, both sides are positive, and the function $x\mapsto x^{1/n_{F}}$ is monotonically increasing on $[0,1]$ for $n_{F}>0$. Therefore, the inequality is preserved under this transformation.
@@
$\Pr(\theta_{ij}\leq t^{1/n_{F}}\mid C_{i},h_{i})=\int_{0}^{r^{1/n_{F}}}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij}$
@@ Appendix C List of Mathematical Notations
$F_{R_{ij}(n_{F})}(t\mid C_{i},h_{i})$ CDF of subdomain $n_{F}$-operation reliability
$F_{R_{i}(n_{F})}(t\mid C_{i},h_{i})$ CDF of domain $n_{F}$-operation reliability
$F_{R_{L}(n_{F})}(t\mid\text{data},\mathcal{H})$ CDF of LLM $n_{F}$-operation reliability